# Conditional Entropies of

### v1

@@ V Extremum Values for $k=1$
V-A The $1$-Deletion Channel
@@ V-A1 Maximum Input Entropy
The next theorem states that the input entropy is maximized by channel outputs with exactly $m$ runs (each of length one).
@@ V-A2 Minimum Input Entropy
The next theorem states that the channel outputs that minimize this entropy have a single run. This extends the results from [6] to the non-binary case.
@@
V-B The $1$-Insertion Channel
@@
VI Average Values for $k=1$
@@
VII The $2$-Deletion Channel
Remark 2.
As highlighted in Lemma 4, the characterization of embedding numbers is slightly different than it was shown in [6]. Thus, finding the accurate value of $\mathsf{min}^{\mathsf{In}}_{2\textrm{-}\mathsf{Del}}(n)$ and the sequences that attain this minimum value becomes more challenging with the corrected characterization due to the involvement of alternating segments. Hence, it is crucial to re-establish the result.
@@ VII The $2$-Deletion Channel
VII-A Minimum Input Entropy
@@
For the ease of exposition, we give the next two lemmas, which we will be useful for deriving the main result of this section, i.e., the value of $\mathsf{min}^{\mathsf{In}}_{2\textrm{-}\mathsf{Del}}(n)$ and the sequences that attain this minimum value.
@@
$+\left(\theta+1\right)\log\left(\theta+1\right)-\theta\log\left(\theta\right)+W_{y_{2}\circ I_{2}({\boldsymbol{y}}_{[3,\ell]})}({\boldsymbol{y}}_{[2,l]}),$
@@
We now analyse the first sum,
$\mathsf{W}_{y_{1}\circ y_{2}\circ I_{2}({\boldsymbol{y}}_{[3,\ell]})}({\boldsymbol{y}})=\sum_{{\boldsymbol{x}}~{}\in~{}y_{2}\circ I_{2}({\boldsymbol{y}}_{[3,\ell]})}\omega_{{\boldsymbol{y}}}(y_{1}\circ{\boldsymbol{x}})\log(\omega_{{\boldsymbol{y}}}(y_{1}\circ{\boldsymbol{x}}))$
@@
Next, let us analyse the second sum,
$\mathsf{W}_{y_{1}\circ\overline{y_{2}}\circ I_{1}({\boldsymbol{y}}_{[2,\ell]})}({\boldsymbol{y}})=\sum_{{\boldsymbol{x}}~{}\in~{}\overline{y_{2}}\circ I_{1}({\boldsymbol{y}}_{[2,\ell]})}\omega_{{\boldsymbol{y}}}(y_{1}\circ{\boldsymbol{x}})\log(\omega_{{\boldsymbol{y}}}(y_{1}\circ{\boldsymbol{x}}))$
@@
$+\left(\theta+1\right)\log\left(\theta+1\right)-\theta\log\left(\theta\right)+W_{y_{2}\circ I_{2}({\boldsymbol{y}}_{[3,\ell]})}({\boldsymbol{y}}_{[2,l]}),$
@@
$+W_{I_{1}({\boldsymbol{y}}_{[2,\ell]})}({\boldsymbol{y}}_{[2,\ell]}).$
Proof.
@@
$\mathsf{W}_{y_{1}\circ\overline{y}_{2}\circ I_{1}({\boldsymbol{y}}_{[2,\ell]})}({\boldsymbol{y}})=\sum_{{\boldsymbol{x}}~{}\in~{}\overline{y}_{2}\circ I_{1}({\boldsymbol{y}}_{[2,\ell]})}\omega_{{\boldsymbol{y}}}(y_{1}\circ{\boldsymbol{x}})\log(\omega_{{\boldsymbol{y}}}(y_{1}\circ{\boldsymbol{x}}))$
@@
$+\left(\theta+1\right)\log\left(\theta+1\right)-\theta\log\left(\theta\right)+W_{y_{2}\circ I_{2}({\boldsymbol{y}}_{[3,\ell]})}({\boldsymbol{y}}_{[2,l]}),$
@@
VIII The $k$-Insertion Channel
In this section, we study the channel outputs ${\boldsymbol{y}}\in\Sigma_{q}^{m}$ that minimize the input entropy of $k\textrm{-}\mathsf{Ins}$ Channel, where $m\triangleq n+k$.
@@ VIII The $k$-Insertion Channel
IX Conclusion
In this work, we analysed the conditional entropies of $k$-deletion and $k$-insertion channels and particularly for $k=1$ and $k=2$, we gave a more elaborate characterization by considering the runs and maximal alternating segments of a sequence. We derived the extremum values and the sequences that attain these extremum values for the input entropy of both $1$-deletion and $1$-insertion channels over arbitrary alphabet size. We also studied the average input entropies for $1$-deletion and $1$-insertion channels. We then derived the minimum values and the sequences that attain these minimum values for the input entropy of $2$-deletion channel and $k$-insertion channel, respectively, over binary alphabet.
@@ IX Conclusion
$+\left(1+\theta\right)\log\left(1+\theta\right)-\left(\theta\right)\log\left(\theta\right)$
$+(m-R+1)(r_{1}+2)\log\left(r_{1}+2\right)-(m-R+1)(r_{1}+1)\log\left(r_{1}+1\right)$
$+\sum_{i=2}^{R}\Big{\{}(r_{1}+2)(r_{i}+1)\log\left((r_{1}+2)(r_{i}+1)\right)\Big{\}}$
$-\sum_{i=2}^{R}\Big{\{}(r_{1}+1)(r_{i}+1)\log\left((r_{1}+1)(r_{i}+1)\right)\Big{\}},$
@@
$+\sum_{i=2}^{R}\Big{\{}(r_{1}+2)(r_{i}+1)\log\left((r_{1}+2)(r_{i}+1)\right)-(r_{1}+1)(r_{i}+1)\log\left((r_{1}+1)(r_{i}+1)\right)\Big{\}}.$
We now assume to the contrary that for the sequence ${\boldsymbol{y}}^{\prime}\in\Sigma_{2}^{m}$ with ${\cal R}{\cal L}({\boldsymbol{y}}^{\prime})\triangleq(r_{1}+1,r_{2},\ldots,r_{j}-1,\ldots,r_{R})$, the following holds,
@@
$>{\cal W}^{\prime}({\boldsymbol{y}}^{\prime})+\binom{r_{1}+4}{2}\log\binom{r_{1}+4}{2}-\binom{r_{1}+3}{2}\log\binom{r_{1}+3}{2}.$
@@
$(m-R+1)(r_{1}+3)\log\left(r_{1}+3\right)-(m-R+1)(r_{1}+2)\log\left(r_{1}+2\right)$
$+\sum_{i=2}^{R}\Big{\{}(r_{1}+3)(r_{i}+1)\log\left((r_{1}+3)(r_{i}+1)\right)-(r_{1}+2)(r_{i}+1)\log\left((r_{1}+2)(r_{i}+1)\right)\Big{\}}$
$-\Big{\{}(r_{1}+3)(r_{j}+1)\log\left((r_{1}+3)(r_{j}+1)\right)-(r_{1}+2)(r_{j}+1)\log\left((r_{1}+2)(r_{j}+1)\right)\Big{\}}$
+$+\Big{\{}(r_{1}+3)(r_{j})\log\left((r_{1}+3)(r_{j})\right)-(r_{1}+2)(r_{j})\log\left((r_{1}+2)(r_{j})\right)\Big{\}}$
$-\Bigg{\{}(m-R+1)(r_{1}+2)\log\left(r_{1}+2\right)-(m-R+1)(r_{1}+1)\log\left(r_{1}+1\right)$
$+\sum_{i=2}^{R}\Big{\{}(r_{1}+2)(r_{i}+1)\log\left((r_{1}+2)(r_{i}+1)\right)-(r_{1}+1)(r_{i}+1)\log\left((r_{1}+1)(r_{i}+1)\right)\Big{\}}\Bigg{\}}.$
After some simplification, it follows that ${\cal W}^{\prime}({\boldsymbol{y}}^{\prime})-{\cal W}^{\prime}({\boldsymbol{y}})$ is
$(m-R+1)\bigg{\{}(r_{1}+3)\log\left(r_{1}+3\right)-2(r_{1}+2)\log\left(r_{1}+2\right)+(r_{1}+1)\log\left(r_{1}+1\right)\bigg{\}}$
$+\sum_{i=2}^{R}(r_{i}+1)\Big{\{}(r_{1}+3)\log\left(r_{1}+3\right)-2(r_{1}+2)\log\left(r_{1}+2\right)+(r_{1}+1)\log\left(r_{1}+1\right)\Big{\}}$
$+(r_{1}+2)\log(r_{1}+2)-(r_{1}+3)\log(r_{1}+3)+(r_{j})\log(r_{j})-(r_{j}+1)\log(r_{j}+1).$
Upon expanding the summation, we get ${\cal W}^{\prime}({\boldsymbol{y}}^{\prime})-{\cal W}^{\prime}({\boldsymbol{y}})$ as
$(r_{1}+2)\log(r_{1}+2)-(r_{1}+3)\log(r_{1}+3)+(r_{j})\log(r_{j})-(r_{j}+1)\log(r_{j}+1)$
$+(2m-r_{1})\bigg{\{}(r_{1}+3)\log\left(r_{1}+3\right)-2(r_{1}+2)\log\left(r_{1}+2\right)+(r_{1}+1)\log\left(r_{1}+1\right)\bigg{\}}$
We know that $(r+1)\log(r+1)-r\log(r)$ is an increasing function w.r.t. $r$. Therefore, the difference ${\cal W}^{\prime}({\boldsymbol{y}}^{\prime})-{\cal W}^{\prime}({\boldsymbol{y}})$ is lower bounded by
$+(r_{1}+2)\log(r_{1}+2)-(r_{1}+3)\log(r_{1}+3)+(m)\log(m)-(m+1)\log(m+1).$
Hence, the difference

### v2

@@ V Extremum and Average Values for $k=1$
V-A Extremum Values for the $1$-Deletion Channel
@@ V-A1 Maximum Input Entropy
The next theorem states that the input entropy of the $1$-deletion channel is maximized by channel outputs with exactly $m$ runs (each of length one).
@@ V-A2 Minimum Input Entropy
The next theorem states that the channel outputs that minimize input entropy of the $1$-deletion channel have a single run. This extends the results from [6] to the non-binary case.
@@
V-B Extremum Values for the $1$-Insertion Channel
@@
The next theorem states that the input entropy of the $1$-insertion channel is maximized by channel outputs with exactly $m$ runs (each of length one).
@@
The next theorem states that the channel outputs that minimize input entropy of the $1$-insertion channel have a single run.
@@
V-C Average Values for the $1$-insertion and $1$-deletion channels
@@
VI Extremum Values for $k\geq 2$
@@
Remark 4.
As highlighted in Lemma 4, the characterization of embedding numbers differs slightly from that presented in [6]. Consequently, determining the exact value of $\mathsf{min}^{\mathsf{In}}_{2\textrm{-}\mathsf{Del}}(n)$, as well as identifying the sequences that attain this minimum, becomes more challenging under the corrected characterization due to the involvement of alternating segments. Accordingly, we re-establish the result in Theorem 9 to reflect the corrected characterization.
VI-A Minimum Input Entropy for $2$-Deletion Channel
@@
For the ease of exposition, we give the next two lemmas, which will be useful for deriving the main result of this section, i.e., the value of $\mathsf{min}^{\mathsf{In}}_{2\textrm{-}\mathsf{Del}}(n)$ and the sequences that attain this minimum value.
@@
$\quad+\left(\theta+1\right)\log\left(\theta+1\right)-\theta\log\left(\theta\right)+W_{y_{2}\circ I_{2}({\boldsymbol{y}}_{[3,\ell]})}({\boldsymbol{y}}_{[2,l]}),$
@@
We now analyse the first term,
$\mathsf{W}_{y_{1}\circ y_{2}\circ I_{2}({\boldsymbol{y}}_{[3,\ell]})}({\boldsymbol{y}})$
$=\sum_{{\boldsymbol{x}}~{}\in~{}y_{2}\circ I_{2}({\boldsymbol{y}}_{[3,\ell]})}\omega_{{\boldsymbol{y}}}(y_{1}\circ{\boldsymbol{x}})\log(\omega_{{\boldsymbol{y}}}(y_{1}\circ{\boldsymbol{x}}))$
@@
Next, we analyse the second term,
$\mathsf{W}_{y_{1}\circ\overline{y_{2}}\circ I_{1}({\boldsymbol{y}}_{[2,\ell]})}({\boldsymbol{y}})$
$=\sum_{{\boldsymbol{x}}~{}\in~{}\overline{y_{2}}\circ I_{1}({\boldsymbol{y}}_{[2,\ell]})}\omega_{{\boldsymbol{y}}}(y_{1}\circ{\boldsymbol{x}})\log(\omega_{{\boldsymbol{y}}}(y_{1}\circ{\boldsymbol{x}}))$
@@
$\quad+\left(\theta+1\right)\log\left(\theta+1\right)-\theta\log\left(\theta\right)+W_{y_{2}\circ I_{2}({\boldsymbol{y}}_{[3,\ell]})}({\boldsymbol{y}}_{[2,l]}),$
@@
$\quad+W_{I_{1}({\boldsymbol{y}}_{[2,\ell]})}({\boldsymbol{y}}_{[2,\ell]}).$
Proof:
@@
$\mathsf{W}_{y_{1}\circ\overline{y}_{2}\circ I_{1}({\boldsymbol{y}}_{[2,\ell]})}({\boldsymbol{y}})$
$=\sum_{{\boldsymbol{x}}~{}\in~{}\overline{y}_{2}\circ I_{1}({\boldsymbol{y}}_{[2,\ell]})}\omega_{{\boldsymbol{y}}}(y_{1}\circ{\boldsymbol{x}})\log(\omega_{{\boldsymbol{y}}}(y_{1}\circ{\boldsymbol{x}}))$
@@
$\quad+\left(\theta+1\right)\log\left(\theta+1\right)-\theta\log\left(\theta\right)+W_{y_{2}\circ I_{2}({\boldsymbol{y}}_{[3,\ell]})}({\boldsymbol{y}}_{[2,l]}),$
@@
VI-B Minimum Input Entropy for the $k$-Insertion Channel
In this section, we study the channel outputs ${\boldsymbol{y}}\in\Sigma_{q}^{m}$ that minimize the input entropy of $k\textrm{-}\mathsf{Ins}$ Channel, where $m\triangleq n+k$ and $k\geq 2$.
@@ VI-B Minimum Input Entropy for the $k$-Insertion Channel
VII Conclusion
In this work, we analyzed the conditional entropies of $k$-deletion and $k$-insertion channels and particularly for $k=1$ and $k=2$, we gave a more elaborate characterization by considering the runs and maximal alternating segments of a sequence. We derived the extremum values and the sequences that attain these extremum values for the input entropy of both $1$-deletion and $1$-insertion channels over arbitrary alphabet size. We also studied the average input entropies for $1$-deletion and $1$-insertion channels. We then derived the minimum values and the sequences that attain these minimum values for the input entropy of $2$-deletion channel and $k$-insertion channel over binary alphabet.
VII-A Future Work and Open Problems
Although this paper presents results of significant importance in characterizing the entropy of $k$-deletion and $k$-insertion channels, several open problems remain for future exploration:
  • Maximal Input Entropy for $k=2$. While we provide a complete characterization of the sequences that minimize input entropy, the precise identification of the sequences that maximize input entropy for the $2$-deletion and $2$-insertion channels is still unresolved. Determining these sequences and the associated maximal entropy values remains an important open direction.
  • Extending Analysis to $k>2$. The extremization and average-case analysis for $k$-deletion and $k$-insertion channels with $k>2$ remain largely open. A key challenge lies in the combinatorial explosion of the number of embeddings and the absence of a tractable description of the conditional probability distribution for general $k$. The structural techniques developed in this paper—based on runs and alternating segments—do not extend readily to higher values of $k$, suggesting the need for new analytical tools.
  • Non-Uniform Input Distributions. All results in this work are obtained under the assumption of uniform transmission. Analyzing extremal and average conditional entropy values under general or capacity-achieving input distributions is an important next step. This line of inquiry is particularly relevant for understanding the channel capacity of $k$-deletion/insertion channels, where the optimal input distribution is currently unknown.
@@ VII-A Future Work and Open Problems
$\quad+\left(1+\theta\right)\log\left(1+\theta\right)-\left(\theta\right)\log\left(\theta\right)$
$\quad+(m-R+1)(r_{1}+2)\log\left(r_{1}+2\right)-(m-R+1)(r_{1}+1)\log\left(r_{1}+1\right)$
$\quad+\sum_{i=2}^{R}\Big{\{}(r_{1}+2)(r_{i}+1)\log\left((r_{1}+2)(r_{i}+1)\right)\Big{\}}$
$\quad-\sum_{i=2}^{R}\Big{\{}(r_{1}+1)(r_{i}+1)\log\left((r_{1}+1)(r_{i}+1)\right)\Big{\}},$
@@
$\quad+\sum_{i=2}^{R}\Big{\{}(r_{1}+2)(r_{i}+1)\log\left((r_{1}+2)(r_{i}+1)\right)-(r_{1}+1)(r_{i}+1)\log\left((r_{1}+1)(r_{i}+1)\right)\Big{\}}.$
We now assume to the contrary that for the sequence ${\boldsymbol{y}}^{\prime}\in\Sigma_{2}^{m}$ with ${\cal R}{\cal L}({\boldsymbol{y}}^{\prime})\triangleq(r_{1}+1,r_{2},\ldots,r_{j}-1,\ldots,r_{R})$, the following holds:
@@
>$\quad$> ${\cal W}^{\prime}({\boldsymbol{y}}^{\prime})+\binom{r_{1}+4}{2}\log\binom{r_{1}+4}{2}-\binom{r_{1}+3}{2}\log\binom{r_{1}+3}{2}.$
@@
${\cal W}^{\prime}({\boldsymbol{y}}^{\prime})-{\cal W}^{\prime}({\boldsymbol{y}})$ $=(m-R+1)(r_{1}+3)\log\left(r_{1}+3\right)-(m-R+1)(r_{1}+2)\log\left(r_{1}+2\right)$
$\quad+\sum_{i=2}^{R}\Big{\{}(r_{1}+3)(r_{i}+1)\log\left((r_{1}+3)(r_{i}+1)\right)-(r_{1}+2)(r_{i}+1)\log\left((r_{1}+2)(r_{i}+1)\right)\Big{\}}$
$\quad-\Big{\{}(r_{1}+3)(r_{j}+1)\log\left((r_{1}+3)(r_{j}+1)\right)-(r_{1}+2)(r_{j}+1)\log\left((r_{1}+2)(r_{j}+1)\right)\Big{\}}$
+$\quad+\Big{\{}(r_{1}+3)(r_{j})\log\left((r_{1}+3)(r_{j})\right)-(r_{1}+2)(r_{j})\log\left((r_{1}+2)(r_{j})\right)\Big{\}}$
$\quad-\Bigg{\{}(m-R+1)(r_{1}+2)\log\left(r_{1}+2\right)-(m-R+1)(r_{1}+1)\log\left(r_{1}+1\right)$
$\quad+\sum_{i=2}^{R}\Big{\{}(r_{1}+2)(r_{i}+1)\log\left((r_{1}+2)(r_{i}+1)\right)-(r_{1}+1)(r_{i}+1)\log\left((r_{1}+1)(r_{i}+1)\right)\Big{\}}\Bigg{\}}.$
After some simplification, it follows that
${\cal W}^{\prime}({\boldsymbol{y}}^{\prime})-{\cal W}^{\prime}({\boldsymbol{y}})$ $=(m-R+1)\bigg{\{}(r_{1}+3)\log\left(r_{1}+3\right)-2(r_{1}+2)\log\left(r_{1}+2\right)+(r_{1}+1)\log\left(r_{1}+1\right)\bigg{\}}$
$\quad+\sum_{i=2}^{R}(r_{i}+1)\Big{\{}(r_{1}+3)\log\left(r_{1}+3\right)-2(r_{1}+2)\log\left(r_{1}+2\right)+(r_{1}+1)\log\left(r_{1}+1\right)\Big{\}}$
$\quad+(r_{1}+2)\log(r_{1}+2)-(r_{1}+3)\log(r_{1}+3)+(r_{j})\log(r_{j})-(r_{j}+1)\log(r_{j}+1).$
Upon expanding the summation, we get
${\cal W}^{\prime}({\boldsymbol{y}}^{\prime})-{\cal W}^{\prime}({\boldsymbol{y}})$ $=(r_{1}+2)\log(r_{1}+2)-(r_{1}+3)\log(r_{1}+3)+(r_{j})\log(r_{j})-(r_{j}+1)\log(r_{j}+1)$
$\quad+(2m-r_{1})\bigg{\{}(r_{1}+3)\log\left(r_{1}+3\right)-2(r_{1}+2)\log\left(r_{1}+2\right)+(r_{1}+1)\log\left(r_{1}+1\right)\bigg{\}}.$
We know that $(r+1)\log(r+1)-r\log(r)$ is an increasing function w.r.t. $r$. Therefore, we get the folloing lower bounded:
${\cal W}^{\prime}({\boldsymbol{y}}^{\prime})-{\cal W}^{\prime}({\boldsymbol{y}})$ $\geq(2m-r_{1})\bigg{\{}(r_{1}+3)\log\left(r_{1}+3\right)-2(r_{1}+2)\log\left(r_{1}+2\right)+(r_{1}+1)\log\left(r_{1}+1\right)\bigg{\}}$
$\quad+(r_{1}+2)\log(r_{1}+2)-(r_{1}+3)\log(r_{1}+3)+(m)\log(m)-(m+1)\log(m+1).$
Hence, the following expression: